# RichRAG: Crafting Rich Responses for Multi-faceted Queries in Retrieval-Augmented Generation

### v2

@@ Appendix A Data Pre-process
Specifically, for the WikiPassageQA dataset, we prompt GPT-4 to generate appropriate and accurate sub-aspects and sub-answers for each data point to support our study, where sub-answers are split from the original long-form answers and sub-aspects are closely related to original questions and sub-answers. The content of the prompt is shown in Prompt E. The WikiAsp dataset is devised for generating aspect-based summaries of entities from 20 domains. Each piece of data is built from a Wikipedia article, consisting of various aspects of this article and aspect-based summaries, each target summary corresponds to an aspect. We follow Jiang et al. (2023) to convert it into open-domain QA settings and introduce some additional operations to make it suitable for our experimental settings. Firstly, samples with more than two aspects are retained and the remaining parts are removed to ensure the multifaceted nature of the experimental samples. Then, due to the expensive costs of experiments and the large amount of the whole dataset, we evenly sample subsets from each domain to build the experimental samples, and split training, validation, and test sets according to the ratio of 10:1:1. Finally, we insert the title of the original Wikipedia article into a template: “Generate a summary about title” to mimic the real question format, hence constructing the question of each sample data. The aspects and aspect-based summaries are treated as sub-aspects of the question and sub-answers. We concatenate these sub-answers to build the long-form answer for each sample. The statistical information of datasets is presented in Table 3.
@@ Appendix B Implementation Details
The sub-aspect explorer is implemented by Llama-2-7B-chat Touvron et al. (2023). We set the learning rate as 5e-5, batch size as 64, and use AdamW Loshchilov and Hutter (2019) to fine-tune it. For the multi-faceted retriever, the number of retrieved documents for each sub-aspect is set as 50. The maximum capacity of the pool is 290 for WikipasssageQA and 270 for WikiAsp. For the generative list-wise ranker, We base on Flan-T5-base Chung et al. (2022) to initialize it and rerank the top-10 final documents as provided external knowledge for the generator. In the SFT stage, we set the learning rate as 5e-5, batch size as 64, temperature $\tau$ as 0.1, and optimize the ranker to generate top-$10$ ranked document IDs with AdamW algorithm. In the RL stage, Llama-2-13B-chat is chosen as the generator, $\mathcal{G}$, providing reward feedback to the policy model. Then, we set $\mu$ as 0.1 to build 6,000 training pairs for the DPO algorithm. The batch size and learning rate are set as 32 and 3e-6 respectively to further optimize our ranker via the DPO strategy. We followed Izacard et al. (2024) to consider the Dec. 20, 2021, Wikipedia dump as our knowledge base and utilize BGE-en-base Xiao et al. (2023) as our fixed retriever. Our experiments are conducted on the platform with four NVIDIA A100-SXM4-80GB GPUs. We will release our codes upon the acceptance of our study.
@@
Appendix C Analysis of LLMs’ Preferences
@@
Appendix D Impact of Number of External Documents
@@
Appendix E Case Study
To validate that our proposed framework has the ability to provide rich and comprehensive responses for multi-faceted questions, we further demonstrate a case study to compare the generated responses of RichRAG and a RAG baseline that directly generates responses based on retrieved documents. In Box E, we provide the input question, its labeled sub-aspects, and predictions from different compared models. The parts of generated responses are highlighted by different colors to indicate the comprehensiveness of different results. According to the visualized results, it is obvious that RichRAG could predict the query’s sub-aspects accurately and provide rich responses that satisfy these potential intents. However, the baseline model only offers a general and incomplete response without in-depth consideration of the various underlying intentions of the user, which may implicitly impact user satisfaction. This result further proves the ability of RichRAG to identify and explicitly model the query’s sub-aspects, thereby providing satisfying ranking results to enhance the richness and comprehensiveness of generated responses.

### v3

@@ Appendix A Data Pre-process
Specifically, for the WikiPassageQA dataset, we prompt GPT-4 to generate appropriate and accurate sub-aspects and sub-answers for each data point to support our study, where sub-answers are split from the original long-form answers and sub-aspects are closely related to original questions and sub-answers. The content of the prompt is shown in Prompt G. To ensure the quality of reformulated WikiPassageQA, the prompt of GPT-4 is decided via the following steps:
(1) Human-Annotated Demonstrations: We first selected five data examples and manually annotated their sub-aspects and sub-answers. These human-annotated examples are used as demonstrations to prompt the GPT-4 to generate satisfied results.
(2) Calibration with Human verification: Then, we randomly sampled 50 examples and used them as calibration examples. Specifically, we ceaselessly adjusted the content of the prompt based on the generation quality of these calibration examples until we thought the generation quality of these examples was satisfied. The satisfaction rate of human evaluation was 94%.
The WikiAsp dataset is devised for generating aspect-based summaries of entities from 20 domains. Each piece of data is built from a Wikipedia article, consisting of various aspects of this article and aspect-based summaries. Each target summary corresponds to an aspect. We follow Jiang et al. (2023) to convert it into open-domain QA settings and introduce some additional operations to make it suitable for our experimental settings. Firstly, samples with more than two aspects are retained, and the remaining parts are removed to ensure the multifaceted nature of the experimental samples. Then, due to the expensive costs of experiments and the large amount of the whole dataset, we evenly sample subsets from each domain to build the experimental samples and split training, validation, and test sets according to the ratio of 10:1:1. Finally, we insert the title of the original Wikipedia article into a template: “Generate a summary about title” to mimic the real question format, hence constructing the question of each sample data. The aspects and aspect-based summaries are treated as sub-aspects of the question and sub-answers. We concatenate these sub-answers to build the long-form answer for each sample.
The statistical information of our datasets is presented in Table 2.
@@ Appendix B Implementation Details
The sub-aspect explorer is implemented by Llama-2-7B-chat Touvron et al. (2023). We set the learning rate as 5e-5, batch size as 64, and use AdamW Loshchilov and Hutter (2019) to fine-tune it. For the multi-faceted retriever, the number of retrieved documents for each sub-aspect is set as 50. The maximum capacity of the pool is 290 for WikipasssageQA and 270 for WikiAsp. For the generative list-wise ranker, We base on Flan-T5-base Chung et al. (2022) to initialize it and rerank the top-10 final documents as provided external knowledge for the generator. In the SFT stage, we set the learning rate as 5e-5, batch size as 64, temperature $\tau$ as 0.1, and optimize the ranker to generate top-$10$ ranked document IDs with AdamW algorithm. In the RL stage, Llama-2-13B-chat is chosen as the generator, $\mathcal{G}$, providing reward feedback to the policy model. Then, we set $\mu$ as 0.1 to build 6,000 training pairs for the DPO algorithm. The batch size and learning rate are set as 32 and 3e-6 respectively to further optimize our ranker via the DPO strategy. We followed Izacard et al. (2024) to consider the Dec. 20, 2021, Wikipedia dump as our knowledge base and utilize BGE-en-base Xiao et al. (2023) as our fixed retriever. All our baselines are optimized and evaluated by the same training, validation, and test datasets. Since the WikiAsp dataset has no relevance labels on documents, for each query,we view the documents whose matching score (the average of rouge scores evaluated by the ground truth answer) is higher than 0.5 as relevant documents to support the training of our baseline ranking models. Our experiments are conducted on the platform with four NVIDIA A100-SXM4-80GB GPUs. We will release our codes upon the acceptance of our study.
Models WikiPassageQA WikiAsp
MAP N1 N3 N5 N10 NCOM MAP N1 N3 N5 N10 NCOM
Retriever .3201 .2476 .2933 .3498 .3951 .7006 .1780 .1003 .1678 .1953 .2294 .3926
RankT5 (point) .5200 .4808 .5176 .5424 .5916 .7865 .3712 .3230 .3704 .3942 .4043 .4739
RankT5 (pair) .3602 .2740 .3367 .3863 .4298 .7395 .3299 .2088 .3274 .3717 .3894 .5297
LDIST (point) .5346 .5024 .5261 .5573 .5979 .7608 .4051 .3668 .4078 .4213 .4280 .5043
LiT5 (list) .3756 .2813 .3661 .3968 .4412 .7465 .2757 .1696 .2742 .3142 .3318 .5442
RAG-Fusion .2632 .1923 .2348 .2754 .3317 .5771 .2040 .1188 .1844 .2271 .2608 .4415
+RankT5 (point) .3962 .3510 .3831 .4238 .4641 .6543 .1591 .0681 .1411 .1742 .2177 .4610
+RankT5 (pair) .3946 .3029 .3669 .4202 .4758 .6717 .3297 .2376 .3249 .3583 .3869 .4992
+LDIST (point) .3746 .3245 .3654 .3951 .4385 .6370 .3345 .2987 .3241 .3505 .3733 .5138
+LiT5 (list) .3114 .2308 .2811 .3244 .3839 .6547 .2737 .1995 .2606 .3014 .3255 .5196
BGM (list) .2614 .2476 .2480 .2626 .2864 .7988 .1382 .0727 .1152 .1497 .1921 .3887
RichRAG (list) .5444 .5240 .5359 .5663 .6035 .8065 .4880 .4556 .4962 .5047 .5067 .8935
Table 3: Overall ranking results. The best and second-best results are in bold and underlined, respectively.
Appendix C Ranking Performance
To prove the effectiveness of our proposed list-wise ranking module, we also evaluate the ranking modules’ performance of our method and baselines. We indicate the types of ranking algorithms, including point-wise and list-wise. We further implement a pairwise version of the RankT5 model by using the following training objective to optimize it:
$\mathcal{L}=\!\!\!\!\!\!\!\!\!\!\!\!\!\!\sum_{d_{1},d_{2}\in\mathcal{D},R(d_{1})>R(d_{2})}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\max(0,s(d_{2})-s(d_{1})+\gamma),\;\;\gamma=1,$
where $\mathcal{D}$ denotes the candidate documents, $R(d)$ represents the relevance label of the document and $s(d)$ is the predicted score of the document.
Specifically, we select some widely-used metrics, MAP and NDCG@k(k=1,3,5,10) to assess the ability to predict the document relevance of models. Furthermore, our model not only focuses on document relevance but also the comprehensiveness of the provided ranking lists. Thus, we propose a novel ranking metric, Normalized Comprehensiveness (NCOM), to assess the comprehensiveness of ranking lists, following the way to build the silver ranking targets (introduced in Section 3.4.1). This metric considers both the document relevance and the coverage of ranking lists on sub-aspects, hence reliably evaluating the comprehensiveness of ranking lists. Its calculation is presented below:
Given a question $q$, its sub-answers $\{a_{1},...a_{n}\}$, a pool of candidate document $\mathcal{P}$, a ranking list generated by a ranking model, $L=[d^{1},...,d^{K}]$, and a silver ranking list, $L^{*}$, generated by Eq. 8 in our paper, we calculate the generated and silver ranking lists’ comprehensiveness scores as follow:
$COM(L)$ $=\sum_{t=1}^{K}\sum_{i=1}^{n}w^{t}_{i}\cdot\phi(d^{t},a_{i}),$ (13)
$w^{t}_{i}$ $=1-\mathrm{Norm}_{i}(\max_{d\in L[:t]}\phi(d,a_{i}))$
where $w^{t}_{i}$ denotes the importance of $i$-th sub-aspects at the $t$-th step (which is the same as the Equation 9 in our paper.), and $\phi(\cdot,\cdot)$ calculates the similarity between two sentences, i.e., average of Rouge-2, and Rouge-L scores.
The final normalized comprehensiveness score is computed by normalizing $COM(L)$ by $COM(L^{*})$:
$NCOM=COM(L)/COM(L^{*}).$
The evaluation results are shown in Table 2. From the experimental results, we can find that our method outperforms all types of ranking algorithms on both datasets and all metrics, especially on NCOM. This phenomenon proves that our ranking model could consider the document’s relevance and list comprehensiveness simultaneously, hence providing comprehensive external knowledge and stimulating LLMs to generate better responses. We also notice that even though LiT5 and BGM are list-wise ranking models, their optimization does not explicitly consider the relationships among documents and, hence cannot provide satisfactory results. It also confirms the superiority of our proposed training algorithm.
@@ Appendix C Ranking Performance
Figure 7: Subset experiments with different sub-aspect amounts.
Appendix D Impact of Sub-aspect Amount
Appendix E Analysis of LLMs’ Preferences
@@
Appendix F Impact of Number of External Documents
@@
Appendix G Case Study
To validate that our proposed framework has the ability to provide rich and comprehensive responses for multi-faceted questions, we further demonstrate a case study to compare the generated responses of RichRAG and an RAG baseline that directly generates responses based on retrieved documents. In Box G, we provide the input question, its labeled sub-aspects, and predictions from different compared models. The parts of generated responses are highlighted by different colors to indicate the comprehensiveness of different results. According to the visualized results, it is obvious that RichRAG could predict the query’s sub-aspects accurately and provide rich responses that satisfy these potential intents. However, the baseline model only offers a general and incomplete response without in-depth consideration of the various underlying intentions of the user, which may implicitly impact user satisfaction. This result further proves the ability of RichRAG to identify and explicitly model the query’s sub-aspects, thereby providing satisfying ranking results to enhance the richness and comprehensiveness of generated responses. This result also reveals that though LLMs have the potential ability to generate comprehensive answers, comprehensive ranking lists that correctly cover the query’s sub-aspects can better stimulate LLMs’ ability to generate comprehensive responses. This can also be verified by our quantitative experiments where our model exhibits the best performance among baselines.